OpenStudy (animalover02):

I WILL METAL!!!! & FAN!!!!!!!!!!!!!:) y is greater than or equal to 1-3x present this on a graph.

563blackghost (563blackghost):

`greater than or equal to` this sign is represented by \(\large\bf{\ge}\) or \(\large\bf{\le}\) ?

OpenStudy (animalover02):

> that way

OpenStudy (ankitshaw):

use y = 1-3x to draw the line and then shade the portion that satisfies y>= 1-3x

OpenStudy (animalover02):

i have 4 answer in graphs but i dont want to draw them all out so ill just have someone draw the right one then pick it out of the answer

563blackghost (563blackghost):

Nice so we have... \(\huge\bf{y \ge 1 - 3x}\) Now you would first plug this as any slope-intercept equation. Due to the fact it is \(\large\bf{\ge}\) the line would be bold (if it were > or < it would be dotted).

563blackghost (563blackghost):

Take a screenshot of it.

OpenStudy (animalover02):

ok give me a min.

OpenStudy (animalover02):

i dont know how to put the screenshot on her though..i can draw the graphs but it might take 5 min. for each graph and there are 4 of them

563blackghost (563blackghost):

To take a screenshot you press `PrtScn` on the corner of your keyboard. Once you do that open your default paint and click on the pasteboard in the corner (or right-click and paste). Once that is done crop it to hide personal information and then save. You would then press `Attach File` and find the photo and then send :)

OpenStudy (animalover02):

sounds comlicated can i just draw them. ill go fast

563blackghost (563blackghost):

Its not that complicated, but you can if you'd like.

563blackghost (563blackghost):

\(\huge\bf{y \ge 1 - 3x}\) Now you would first plug this as any slope-intercept equation. Due to the fact it is \(\large\bf{\ge}\) the line would be bold (if it were > or < it would be dotted). Due to the fact `y` is greater than the value of y will increase. Due to this the shaded region will go to the right of the graph.

OpenStudy (animalover02):